A kicked football leaves the ground at an angle θ = 45with the horizontal has an initial speed of 25 m/s. Determine the distance of X. Acceleration due to gravity is 9.8 m/s2.



1
Expert's answer
2022-05-19T05:42:42-0400
"R=\frac{v_0^2\sin2\theta}{g}=\frac{25^2\sin90^\circ}{9.8}=64\:\rm m"